# Enumerative Geometry of Quantum Periods

### v1

@@ 5.5. Higher genus open-log correspondence
From previously shown results in Section 5, we have a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ that relates the generating function of two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to the genus-$g$, winding-1, 1-boundary component, open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ in framing-0,
@@
Assume that $\mathbb{P}$ is toric, and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up. Then, we have,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ \sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\left[(-1)^{\beta\cdot E+g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta$
@@
Thus, by Theorem 5.8, we can express log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ by open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$, and vice versa.
@@
Open-log correspondences in all genera have been proven when the anti-canonical divisor is singular. In [BBvG], an open-log correspondence is proven for Looijenga pairs satisfying a deformation property, and is extended to quasi-tame Looijenga pairs in [BrSc]. In [Sch], an all genus open-log correspondence is established for two-component Looijenga pairs.
Example 5.10.
Let $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. We may write $O_{1}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))=n_{1}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))-\frac{1}{24}n_{0}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))$ by the open multiple cover formulas [MV]. Specializing Theorem 5.8 to genus-1 and degree-4, we have the following equality,
$O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))=\frac{1}{11}R_{1,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)+\frac{11}{24}R_{0,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)\\ -3R_{0,(3)}(\mathbb{F}_{1}(\log\pi^{*}E-C),\pi^{*}H)-2R_{0,(1)}(\mathbb{F}_{1}(\log\pi^{*}E-C),C)R_{0,(2)}(\mathbb{F}_{1}(\log\pi^{*}E-C),F)$
We can compute that,
$\Delta(1,4H)=O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))-\frac{1}{11}R_{1,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)+\frac{1}{24}n_{0}(K_{\mathbb{F}_{1}},\pi^{*}4H-C)$
where $\Delta(1,4H)$ is defined in Equation 5.2.
The open Gromov-Witten invariant $O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))$ was computed by localization to be $\frac{-3313}{12}$ [GZ]. The relevant right hand side invariants $R_{g,(11,1)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)$ and $R_{0}(\mathbb{F}_{1}(\log\pi^{*}E-C),\beta)$ can be computed by $q$-refined tropical curve counting in the scattering diagram of $(\mathbb{C}\mathbb{P}^{2},E)$ ([GRZ], Appendix B) and/or using the formula of [vGGR]. The first equality above is explicitly,
$\frac{-3313}{12}=\frac{1}{11}\cdot(-18513)+\frac{11}{24}\cdot 3146-3\cdot 9-2\cdot 1\cdot 4.$
@@
Recall from Theorem 4.2 that $\vartheta_{1}(q)$ or equivalently $M(Q,q)$ encode the higher-genus, two-pointed log invariants of $(\mathbb{P},E)$. We note that it is not the generating series of higher-genus, winding-1, 1 boundary component, open invariants of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$. As an example, consider $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. We have from Example 2.27 that,
@@
whereas the higher-genus open invariants of $L\subset K_{\mathbb{C}\mathbb{P}^{2}}$ can be computed by the Topological Vertex (Equation (9.10), [AKMV]),
@@
where $z=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$. The coefficient of $z^{g}Q^{d}$ in $\widehat{f}_{\square\,\varnothing\,\varnothing}$ is the negative of the genus-$g$, degree-$d$, winding-1, 1 boundary component, LMOV invariant of an outer AV-brane in local $\mathbb{C}\mathbb{P}^{2}.$ For an outer AV-brane, the winding-1, LMOV invariant is equal to the winding-1, open-BPS invariant $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$.  [^6] In general, LMOV invariants are related to open-BPS invariants by a linear transformation defined by characters of the symmetric group [AKMV]. Note that in the limit as $z\rightarrow 0$ or $q\rightarrow 1$, we recover the open mirror map $M(Q,q)$ from $\widehat{f}_{\square\,\varnothing\,\varnothing}$. From Example 2.27, we see that $\widehat{f}_{\square\,\varnothing\,\varnothing}\neq M(Q,q)$, with the difference $M(Q,q)-\widehat{f}_{\square\,\varnothing\,\varnothing}$ effectively described by Theorem 5.8.
@@
We provide a computational check of Corollary 5.4 and Theorem 5.8 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$ for curve classes $\beta=dH$ for $d\leq 4$ and all genera. To perform calculations with the topological vertex, we use the diagrams in Figure 5.2. The winding-1, open topological string partition function $Z_{K_{\mathbb{C}\mathbb{P}^{2}}}(V)$ of an outer $AV$-brane in local $\mathbb{P}^{2}$ in framing-0 is given by,
$Z_{\mathbb{C}\mathbb{P}^{2}}(V)=\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}(-1)^{\sum_{i}|\lambda_{i}|}Q^{\left(\sum_{i}|\lambda_{i}|\right)H}q^{\sum_{i}\kappa_{\lambda_{i}}}C_{\square\lambda_{2}\lambda^{T}_{3}}C_{\varnothing\lambda_{1}\lambda_{2}^{T}}C_{\varnothing\lambda_{3}\lambda_{1}^{T}}Tr_{\square}V,$
where the topological vertex is encoded by $C_{\lambda_{1},\lambda_{2},\lambda_{3}},$ a Laurent series in $q^{\frac{1}{2}}$ with integer coefficients indexed by three partitions. The closed topological string partition function $Z_{\mathbb{F}_{1}}$ of local $\mathbb{F}_{1}$ with exceptional curve $C$ is given by,
@@
Figure 5.2. The toric diagrams of local $\mathbb{C}\mathbb{P}^{2}$ (left) and local $\mathbb{F}_{1}$ (right). Internal edges have representations $R_{i}$ attached, where for the open calculation of local $\mathbb{C}\mathbb{P}^{2}$ with winding $1$, we have the fundamental representation $\square$ on the external edge.
We begin with degree one, $d=1$. Then all representation labels are either trivial or $\square$, simplifying the expressions for the vertex functions $C_{\lambda\mu\nu}$ (see pg.53, [AKMV] for explicit calculations). Extracting integer invariants, we get
@@
(5.14) $n^{open}_{0}(K_{\mathbb{C}\mathbb{P}^{2}}/L,H+\beta_{0},(1))=2,\quad n^{open}_{g}(K_{\mathbb{C}\mathbb{P}^{2}}/L,H+\beta_{0},(1))=0\text{ for }g>0$
For degrees $2\leq d\leq 4$, we can use the calculations for the refined topological vertex in [IKV], Section 5.6.1, then un-refine them. These are listed in 1.
@@ 3 $(\frac{1}{2},4)\oplus(0,\frac{7}{2})\oplus(0,\frac{5}{2})$ $-(32+9z)$
The table lists the Gopakumar-Vafa invariants of local $\mathbb{F}_{1}$ in curve classes $(d-1)B+dF$, with the $SU(2)_{L}\times SU(2)_{R}$ spin content of BPS states supported on those curves. To un-refine, we want to ignore $j_{R},$ meaning each such representation contributes a factor of $(2j_{R}+1),$ so $(j_{L},j_{R})$ generates a term $(2j_{R}+1)(q^{-2j_{L}}+q^{-2j_{L}+2}+\ldots+q^{2j_{L}-2}+q^{2j_{L}})$ to the curve counting function (see also [HKR], Appendix B for some of these BPS invariants). For given $d$, we compute the total contribution from all of the $(j_{L},j_{R})$. After making a change of variables $z:=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$, the computed expression matches, up to a convention-dependent overall sign, the coefficient of $Q^{d}$ in $-\widehat{f}_{\square\,\varnothing\,\varnothing}$ (defined in Section 5.1) containing the open-BPS invariants in curve class $dH+\beta_{0}$. [^7] In Table 1, the coefficient of $(0,4)$ is listed in [IKV] as 1 instead of 2 — a typo, we believe. Indeed, the coefficient should be 2 in order to recover the genus-0 BPS invariant of local $\mathbb{F}_{1}$ in class $3B+4F$ in the limit $q\rightarrow 1$, which is 286. This invariant is computed in [LLW] [CKYZ] [AKMV]. This completes the verification of Corollary 5.4 for $\mathbb{C}\mathbb{P}^{2}$ for degrees $d\leq 4.$
Appendix A Combinatorial results
A.1. Laurent series inversion and substitution
@@
Lemma A.1 (Lagrange inversion for Laurent series).
@@
Lemma A.4 gives $\operatorname{Res}_{u}(x^{\prime}(u)/x(u))=\operatorname{ord}(x)$ and $\operatorname{Res}_{u}(\partial_{u}\widetilde{f}(x(u))=0$, since $\operatorname{im}(\partial_{u})=\operatorname{ker}(\operatorname{Res}_{u})$. Hence, $\operatorname{Res}_{u}\left(f(x(u))x^{\prime}(u)\right)=\operatorname{ord}(x)\operatorname{Res}_{x}\left(f(x)\right)$ as claimed. The second statement follows from the first statement applied to $f(x)/x$. ∎
A.2. Bell polynomial identities

### v2

@@ 5.4. Higher genus open-log correspondence
From results in Sections 5.2 and 5.3, we have a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ that relates the generating function of two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to the genus-$g$, winding-1, 1-boundary component, open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ in framing-0,
@@
Let $\mathbb{P}$ be a toric Fano surface with smooth anticanonical divisor $E$. Let $L$ be an outer Aganagic-Vafa brane in $K_{\mathbb{P}}$. Then, we have,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ \sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\left[(-1)^{\beta\cdot E+g}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta$
@@
Thus, by Theorem 5.13, we can express log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ by open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$, or equivalently the open Gromov-Witten invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$, and vice versa.
Example 5.14.
Let $\mathbb{P}=\mathbb{P}^{2}$. We write $O_{1}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))=-n_{1}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))+\frac{1}{24}n_{0}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))$ by the open multiple cover formulas [MV]  [^8] In order to match with open Gromov-Witten invariants computed in [AKMV] [LLW] [GZ], we introduce a global sign to the open multiple cover formula in [MV], Equation 2.10.. Specializing Theorem 5.13 to genus-1 and degree-4, we have,
$O_{1}(K_{\mathbb{P}^{2}}/L,4H+\beta_{0},(1))=\frac{1}{11}R_{1,(1,11)}(\mathbb{P}^{2}(\log E),4H)+\frac{11}{24}R_{0,(1,11)}(\mathbb{P}^{2}(\log E),4H)\\ -3R_{0,(3)}(\mathbb{F}_{1}(\log\pi^{*}E-C),\pi^{*}H)-2R_{0,(1)}(\mathbb{F}_{1}(\log\pi^{*}E-C),C)R_{0,(2)}(\mathbb{F}_{1}(\log\pi^{*}E-C),F)$
From Equation 5.2, we can compute that,
$\Delta(1,4H)=O_{1}(K_{\mathbb{P}^{2}}/L,4H+\beta_{0},(1))-\frac{1}{11}R_{1,(1,11)}(\mathbb{P}^{2}(\log E),4H)+\frac{1}{24}n_{0}(K_{\mathbb{F}_{1}},\pi^{*}4H-C)$
where $n_{0}(K_{\mathbb{F}_{1}},\pi^{*}4H-C)=11R_{0,(11,1)}(\mathbb{P}^{2}(\log E),4H)$ by the log-local principle [vGGR] and [GRZ], Corollary 6.6.
The open Gromov-Witten invariant $O_{1}(K_{\mathbb{P}^{2}}/L,4H+\beta_{0},(1))$ was computed by localization to be $\frac{-3313}{12}$ [GZ]. The relevant right hand side invariants $R_{g,(11,1)}(\mathbb{P}^{2}(\log E),4H)$ and $R_{0}(\mathbb{F}_{1}(\log\pi^{*}E-C),\beta)$ can be computed by $q$-refined tropical curve counting in the scattering diagram of $(\mathbb{P}^{2},E)$ ([GRZ], Appendix B) and/or using the formula of [vGGR]. The first equality above is explicitly,
$\frac{-3313}{12}=\frac{1}{11}\cdot(-18513)+\frac{11}{24}\cdot 3146-3\cdot 9-2\cdot 1\cdot 4$
@@
Open-log correspondences in all-genus have been proven when the anti-canonical divisor is singular. In [BBvG], an open-log correspondence is proven for Looijenga pairs satisfying a deformation property, and is extended to quasi-tame Looijenga pairs in [BrSc]. In [Sch], an all genus open-log correspondence is established for two-component Looijenga pairs.
@@
Recall from Theorem 4.2 that $\vartheta_{1}(q)$ or equivalently $M(Q,q)$ encode the higher-genus, two-pointed log invariants of $(\mathbb{P},E)$. We note that, unlike in genus-0, $M(Q,q)$ is not the generating series of higher-genus, 1 boundary component, winding-1 open invariants of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$. For $\mathbb{P}=\mathbb{P}^{2}$, we have from Example 2.28 that,
@@
whereas the higher-genus open invariants of $L\subset K_{\mathbb{P}^{2}}$ can be computed by the Topological Vertex ([AKMV], Equation 9.10),
@@
where $z=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$. The coefficient of $z^{g}Q^{d}$ in $\widehat{f}_{\square\,\varnothing\,\varnothing}$ is the genus-$g$, degree-$d$, LMOV invariant of $L\subset K_{\mathbb{P}^{2}}$ in representation $\square$. By Equation 5.8, the LMOV invariant in representation $\square$ is equal to the winding-1, open-BPS invariant $n_{g}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))$. In the limit as $z\rightarrow 0$ or $q\rightarrow 1$, we recover the genus-0, open mirror map $M(Q,q)$ from $\widehat{f}_{\square\,\varnothing\,\varnothing}$. Hence, we see that $\widehat{f}_{\square\,\varnothing\,\varnothing}\neq M(Q,q)$, with the difference $M(Q,q)-\widehat{f}_{\square\,\varnothing\,\varnothing}$ described in Theorem 5.13.
@@
We provide a computational check of Corollary 5.5 and Theorem 5.13 for $\mathbb{P}=\mathbb{P}^{2}$ for curve classes $\beta=dH$ for $d\leq 4$ and all-genus. To perform calculations with the topological vertex, we use the diagrams in Figure 5.2. The winding-1, open topological string partition function $Z_{\mathbb{P}^{2}}(V)$ of an outer $AV$-brane in local $\mathbb{P}^{2}$ in framing-0 is given by,
$Z_{\mathbb{P}^{2}}(V)=\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}(-1)^{\sum_{i}|\lambda_{i}|}Q^{\left(\sum_{i}|\lambda_{i}|\right)H}q^{\sum_{i}\kappa_{\lambda_{i}}}C_{\square\lambda_{2}\lambda^{T}_{3}}C_{\varnothing\lambda_{1}\lambda_{2}^{T}}C_{\varnothing\lambda_{3}\lambda_{1}^{T}}Tr_{\square}V,$
where the topological vertex is given by $C_{\lambda_{1},\lambda_{2},\lambda_{3}},$ a Laurent series in $q^{\frac{1}{2}}$ with integer coefficients indexed by three partitions. The closed topological string partition function $Z_{\mathbb{F}_{1}}$ of local $\mathbb{F}_{1}$ with exceptional curve $C$ is given by,
@@
Figure 5.2. The toric diagrams of local $\mathbb{P}^{2}$ (left) and local $\mathbb{F}_{1}$ (right). Internal edges have representations $R_{i}$ attached, where for the open calculation of local $\mathbb{P}^{2}$ with winding $1$, we have the fundamental representation $\square$ on the external edge.
We begin with degree one, $d=1$. Then all representation labels are either trivial or $\square$, simplifying the expressions for the vertex functions $C_{\lambda\mu\nu}$ (see [AKMV], pg. 53 for explicit calculations). Extracting integer invariants, we get
@@
(5.14) $n^{open}_{0}(K_{\mathbb{P}^{2}}/L,H+\beta_{0},(1))=2,\quad n^{open}_{g}(K_{\mathbb{P}^{2}}/L,H+\beta_{0},(1))=0\text{ for }g>0$
For degrees $2\leq d\leq 4$, we can use the calculations for the refined Topological Vertex in [IKV], §5.6.1, then un-refine them. These are listed in 1.
@@ 3 $(\frac{1}{2},4)\oplus(0,\frac{7}{2})\oplus(0,\frac{5}{2})$ $-(32+9z)$
The table lists the Gopakumar-Vafa invariants of local $\mathbb{F}_{1}$ in curve classes $(d-1)B+dF$, with the $SU(2)_{L}\times SU(2)_{R}$ spin content of BPS states supported on those curves. To un-refine, we want to ignore $j_{R},$ meaning each such representation contributes a factor of $(2j_{R}+1),$ so $(j_{L},j_{R})$ generates a term $(2j_{R}+1)(q^{-2j_{L}}+q^{-2j_{L}+2}+\ldots+q^{2j_{L}-2}+q^{2j_{L}})$ to the curve counting function (see also [HKR], Appendix B for some of these BPS invariants). For given $d$, we compute the total contribution from all of the $(j_{L},j_{R})$. After making a change of variables $z:=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$, the computed expression matches, up to a convention-dependent overall sign, the coefficient of $Q^{d}$ in $-\widehat{f}_{\square\,\varnothing\,\varnothing}$ (defined in Section 5.1) containing the open-BPS invariants in curve class $dH+\beta_{0}$. [^9] In Table 1, the coefficient of $(0,4)$ is listed in [IKV] as 1 instead of 2 — a typo, we believe. Indeed, the coefficient should be 2 in order to recover the genus-0 BPS invariant of local $\mathbb{F}_{1}$ in class $3B+4F$ in the limit $q\rightarrow 1$, which is 286. This invariant is computed in [LLW] [CKYZ] [AKMV]. This completes the verification of Corollary 5.5 for $\mathbb{P}^{2}$ for degrees $d\leq 4.$
Appendix A Evidence for Conjecture 5.8: GV and LMOV invariants
To provide evidence for Conjecture 5.8, we list here some tables of Gopakumar-Vafa (GV) and Labastida-Mariño-Ooguri-Vafa (LMOV) invariants. The GV invariants $n_{g}(K_{\mathbb{F}_{1}},\pi^{*}\beta-wC)$ for $K_{\mathbb{F}_{1}}$ in curve class $\pi^{*}dH-wC$ come from [HKR]. We abbreviate them $n_{g}(d,w)$ here, for simplicity. The LMOV invariants $N^{\mathrm{LMOV}}_{g,(w)}(K_{\mathbb{P}}/L,\beta)$ in representation $(w)$ are obtained from [AKMV]. [^10] We thank Albrecht Klemm and Jie Gu for calculating $N^{\mathrm{LMOV}}_{g,(w)}(K_{\mathbb{P}^{2}}/L,dH)$ for $w=3,4$ and $d\leq 5$, beyond what appears in [AKMV]. We abbreviate them $N^{\mathrm{LMOV}}_{g,(w)}(d)$ here for simplicity. These are listed for total winding $w\leq 4,$ genus $g\leq 3,$ and degree $d\leq 5.$
$d$ $g=0$ $g=1$ $g=2$ $g=3$
1 -2 0 0 0
2 5 0 0 0
3 -32 9 0 0
4 286 -288 108 -14
5 -3038 6984 -7506 4519
(a) $w=1$: $n_{g}(d,1)=(-1)^{g}N^{\mathrm{LMOV}}_{g,(1)}(d)$
$d$ $g=0$ $g=1$ $g=2$ $g=3$
1 0 0 0 0
2 0 0 0 0
3 7 0 0 0
4 -110 68 -12 0
5 1651 -2938 2353 -992
(b) $w=2$: $n_{g}(d,2)=(-1)^{g}N^{\mathrm{LMOV}}_{g,(2)}(d)$
$d$ $g=0$ $g=1$ $g=2$ $g=3$
1 0 0 0 0
2 0 0 0 0
3 0 0 0 0
4 9 0 0 0
5 -288 300 -116 15
(c) $w=3$: $n_{g}(d,3)=(-1)^{g}N^{\mathrm{LMOV}}_{g,(3)}(d)$
$d$ $g=0$ $g=1$ $g=2$ $g=3$
1 0 0 0 0
2 0 0 0 0
3 0 0 0 0
4 0 0 0 0
5 11 0 0 0
(d) $w=4$: $n_{g}(d,4)=(-1)^{g}N^{\mathrm{LMOV}}_{g,(4)}(d)$
Figure A.1. GV invariants $n_{g}(K_{\mathbb{F}_{1}},\pi^{*}dH-wC)$ equal $(-1)^{g}$ times LMOV invariants $N^{\mathrm{LMOV}}_{g,(w)}(K_{\mathbb{P}^{2}}/L,dH)$ for various genera and degrees, for representations/total winding $w=1,2,3,4$. Tables shown are GV invariants.
Appendix B Combinatorial results
B.1. Laurent series inversion and substitution
@@
Lemma B.1 (Lagrange inversion for Laurent series).
@@
Lemma B.4 gives $\operatorname{Res}_{u}(x^{\prime}(u)/x(u))=\operatorname{ord}(x)$ and $\operatorname{Res}_{u}(\partial_{u}\widetilde{f}(x(u))=0$, since $\operatorname{im}(\partial_{u})=\operatorname{ker}(\operatorname{Res}_{u})$. Hence, $\operatorname{Res}_{u}\left(f(x(u))x^{\prime}(u)\right)=\operatorname{ord}(x)\operatorname{Res}_{x}\left(f(x)\right)$ as claimed. The second statement follows from the first statement applied to $f(x)/x$. ∎
B.2. Bell polynomial identities